# Weyl groups and rigidity of von Neumann algebras

### v1

@@ 1. Introduction and statement of the main results
Since the inclusion $M\subset\mathscr{M}$ is irreducible (see Proposition 2.7 below), we have $\operatorname{Inn}_{M}(\mathscr{M})=\left\{\operatorname{Ad}(u)\mid u\in\mathscr{U}(M^{\prime}\cap\mathscr{M})\right\}=\{\operatorname{id}_{\mathscr{M}}\}$. This further implies that $\mathscr{W}_{G}\cong\operatorname{Aut}_{M}(\mathscr{M})=\operatorname{Out}_{M}(\mathscr{M})$.
The striking feature of Theorem A is that even though the inclusion $M=\operatorname{L}(\Lambda)\subset\operatorname{L}(\Lambda\curvearrowright G/H)=\mathscr{M}$ tends to forget the nonsingular action $\Lambda\curvearrowright G/H$, it still retains the Weyl group $\mathscr{W}_{G}$ via the group isomorphism $\mathscr{W}_{G}\cong\operatorname{Aut}_{M}(\mathscr{M})$. This result can be related to Boutonnet–Houdayer’s noncommutative factor theorem [BH22] (see also [Ho21]). In that respect, set $\mathscr{B}=\operatorname{L}(\Lambda\curvearrowright G/P)$ and naturally regard $M\subset\mathscr{B}\subset\mathscr{M}$. Using the same setting as above and under the higher rank assumption $\sum_{i=1}^{d}\operatorname{rk}_{k_{i}}(\mathbf{G}_{i})\geq 2$, it was shown in [BH22, Theorem B] that the inclusion $M\subset\mathscr{B}$ retains the poset of all intermediate parabolic subgroups. More precisely, it was showed that intermediate von Neumann subalgebras $M\subset\mathscr{N}\subset\mathscr{B}$ are in one-to-one correspondence with intermediate parabolic subgroups $P<Q<G$ via the identification $\mathscr{N}=\operatorname{L}(\Lambda\curvearrowright G/Q)$. We point out that even though the Weyl group $\mathscr{W}_{G}$ can be used to recover the poset of intermediate parabolic subgroups $P<Q<G$, Theorem A is not sufficient in general to recover [BH22, Theorem B]. Indeed, there is no higher rank assumption in Theorem A while the higher rank assumption is essential in the proof of [BH22, Theorem B] because one uses the noncommutative Nevo–Zimmer theorem [BH19, BBH21] in the simple case ($d=1$) and the dichotomy result for equivariant ucp maps [BBHP20] in the product case ($d\geq 2$).
@@ Acknowledgments
CH is grateful to Amine Marrakchi for insightful conversations that led to Proposition 2.7 and to François Thilmany for thought-provoking discussions on the structure and classification of simple algebraic groups.
@@ 3.1. Proof of Theorem A
Using convexity and taking limits and since $c_{w}=0$ for every $w\in\mathscr{W}_{G}$, we obtain $\operatorname{E}(c)=\alpha_{e}p\neq 0$ and so $c\neq 0$. Write $c=v|c|$ for the polar decomposition of $c\in\mathscr{M}$. Then $v\neq 0$ and $\Theta(u)v=vu$ for every $u\in\mathscr{U}(\mathscr{A})$. Therefore, we have $\Theta(a)v=va$ for every $a\in\mathscr{A}$. Since $\mathscr{A}^{\prime}\cap\mathscr{M}=\mathscr{A}$, it follows that $v^{*}v\in\mathscr{A}^{\prime}\cap\mathscr{M}=\mathscr{A}$ and $vv^{*}\in\Theta(\mathscr{A})^{\prime}\cap\mathscr{M}=\Theta(\mathscr{A})$. Moreover, we have $\Theta(v^{*}v)v=v$ which implies that $\Theta(v^{*}v)vv^{*}=vv^{*}$ and so $vv^{*}\leq\Theta(v^{*}v)$. Let $r\in\mathscr{A}$ be the unique projection such that $vv^{*}=\Theta(r)$. Then we have $\Theta(r)v=vr$ which implies that $v^{*}v=v^{*}vr$ and so $v^{*}v\leq r$. This shows that $\Theta(v^{*}v)\leq\Theta(r)=vv^{*}$ and so $\Theta(v^{*}v)=vv^{*}$. Upon cutting down $v$ on the right hand side by a nonzero finite trace projection in $\mathscr{A}$, we may further assume that $v\neq 0$ and $\operatorname{Tr}(v^{*}v)=\operatorname{Tr}(vv^{*})<+\infty$. Since $\Theta|_{\operatorname{L}(\Lambda)}=\operatorname{id}_{\operatorname{L}(\Lambda)}$, we have that $\Theta(a)\sigma_{\gamma}(v)=\sigma_{\gamma}(v)a$ for every $a\in\mathscr{A}$ and every $\gamma\in\Lambda$.

### v2

@@ 1. Introduction and statement of the main results
Since the inclusion $M\subset\mathscr{M}$ is irreducible (see Proposition 2.7 below), we have $\operatorname{Inn}_{M}(\mathscr{M})=\left\{\operatorname{Ad}(u)\mid u\in\mathscr{U}(M^{\prime}\cap\mathscr{M})\right\}=\{\operatorname{id}_{\mathscr{M}}\}$. This further implies that $\mathscr{W}_{G}\cong\operatorname{Aut}_{M}(\mathscr{M})=\operatorname{Out}_{M}(\mathscr{M})$. The striking feature of Theorem A is that even though the inclusion $M=\operatorname{L}(\Lambda)\subset\operatorname{L}(\Lambda\curvearrowright G/H)=\mathscr{M}$ tends to forget the nonsingular action $\Lambda\curvearrowright G/H$, it still retains the Weyl group $\mathscr{W}_{G}$ via the group isomorphism $\mathscr{W}_{G}\cong\operatorname{Aut}_{M}(\mathscr{M})$.
This result is related to Boutonnet–Houdayer’s noncommutative factor theorem [BH22] (see also [Ho21]). In that respect, set $\mathscr{B}=\operatorname{L}(\Lambda\curvearrowright G/P)$ and naturally regard $M\subset\mathscr{B}\subset\mathscr{M}$. Using the same setting as above and under the higher rank assumption $\sum_{i=1}^{d}\operatorname{rk}_{k_{i}}(\mathbf{G}_{i})\geq 2$, it was showed in [BH22, Theorem B] that intermediate von Neumann subalgebras $M\subset\mathscr{N}\subset\mathscr{B}$ are in one-to-one correspondence with intermediate parabolic subgroups $P<Q<G$ via the identification $\mathscr{N}=\operatorname{L}(\Lambda\curvearrowright G/Q)$. In particular, the inclusion $M\subset\mathscr{B}$ retains the abstract poset of all intermediate parabolic subgroups and the rank $\sum_{i=1}^{d}\operatorname{rk}_{k_{i}}(\mathbf{G}_{i})$. However, it is open whether the inclusion $M\subset\mathscr{B}$ retains the Weyl group $\mathscr{W}_{G}$.
Theorem A above implies that the larger inclusion $M\subset\mathscr{M}$ retains the Weyl group $\mathscr{W}_{G}$ via the group isomorphism $\operatorname{Aut}_{M}(\mathscr{M})\cong\mathscr{W}_{G}$ and so in particular the abstract poset of intermediate parabolic subgroups $P<Q<G$. However, Theorem A and [BH22, Theorem B] are of different nature, their proofs do not rely on the same tools and neither theorem implies the other. We emphasize that there is no higher rank assumption in Theorem A while the higher rank assumption is essential in the proof of [BH22, Theorem B] because one uses the noncommutative Nevo–Zimmer theorem [BH19, BBH21] in the simple case ($d=1$) and the dichotomy result for equivariant ucp maps [BBHP20] in the product case ($d\geq 2$).
@@ Acknowledgments
CH is grateful to Amine Marrakchi for insightful conversations that led to Proposition 2.7 and to François Thilmany for thought-provoking discussions on the structure and classification of simple algebraic groups. The authors also thank Uri Bader and Itamar Vigdorovich for their valuable comments.
@@ 3.1. Proof of Theorem A
Using convexity and taking limits and since $c_{w}=0$ for every $w\in\mathscr{W}_{G}\setminus\{e\}$, we obtain $\operatorname{E}(c)=\alpha_{e}p\neq 0$ and so $c\neq 0$. Write $c=v|c|$ for the polar decomposition of $c\in\mathscr{M}$. Then $v\neq 0$ and $\Theta(u)v=vu$ for every $u\in\mathscr{U}(\mathscr{A})$. Therefore, we have $\Theta(a)v=va$ for every $a\in\mathscr{A}$. Since $\mathscr{A}^{\prime}\cap\mathscr{M}=\mathscr{A}$, it follows that $v^{*}v\in\mathscr{A}^{\prime}\cap\mathscr{M}=\mathscr{A}$ and $vv^{*}\in\Theta(\mathscr{A})^{\prime}\cap\mathscr{M}=\Theta(\mathscr{A})$. Moreover, we have $\Theta(v^{*}v)v=v$ which implies that $\Theta(v^{*}v)vv^{*}=vv^{*}$ and so $vv^{*}\leq\Theta(v^{*}v)$. Let $r\in\mathscr{A}$ be the unique projection such that $vv^{*}=\Theta(r)$. Then we have $\Theta(r)v=vr$ which implies that $v^{*}v=v^{*}vr$ and so $v^{*}v\leq r$. This shows that $\Theta(v^{*}v)\leq\Theta(r)=vv^{*}$ and so $\Theta(v^{*}v)=vv^{*}$. Upon cutting down $v$ on the right hand side by a nonzero finite trace projection in $\mathscr{A}$, we may further assume that $v\neq 0$ and $\operatorname{Tr}(v^{*}v)=\operatorname{Tr}(vv^{*})<+\infty$. Since $\Theta|_{\operatorname{L}(\Lambda)}=\operatorname{id}_{\operatorname{L}(\Lambda)}$, we have that $\Theta(a)\sigma_{\gamma}(v)=\sigma_{\gamma}(v)a$ for every $a\in\mathscr{A}$ and every $\gamma\in\Lambda$.
@@ 3.3. Further rigidity results
Remark 3.2.
We point out that Theorem A holds more generally when for every $i\in\{1,\dots,d\}$, $\mathbf{H}_{i}=\mathscr{Z}_{\mathbf{G}_{i}}(\mathbf{S}_{i})<\mathbf{G}_{i}$ is replaced by any Zariski connected $k_{i}$-subgroup $\mathbf{L}_{i}<\mathbf{G}_{i}$ for which $\mathscr{N}_{\mathbf{G}_{i}}(\mathbf{L}_{i})/\mathbf{L}_{i}$ is finite, every coset of $\mathscr{N}_{\mathbf{G}_{i}}(\mathbf{L}_{i})/\mathbf{L}_{i}$ is represented by an element rational over $k_{i}$ and $L=\prod_{i=1}^{d}\mathbf{L}_{i}(k_{i})$ is noncompact.